OpenStudy (mpj4):

How to obtain parametric equation(s) of x^2+y^2-4x=0?

ganeshie8 (ganeshie8):

HINT : complete the square

OpenStudy (mpj4):

(x-2)^2+y^2=4?

OpenStudy (mpj4):

I do not know what kind of answer I should be getting though

OpenStudy (welshfella):

parametris equations are like x = t^2 , y = 2t are third variable is introduced

OpenStudy (mpj4):

So I have to show that this is a circle using that third variable?

ganeshie8 (ganeshie8):

\(x-2 ~=~ 2\cos t\) \(y ~= ~2\sin t\)

ganeshie8 (ganeshie8):

\(x~ =~2+2\cos t\) \(y~=~2\sin t\)

OpenStudy (mpj4):

Would it be ok to ask on how that happened?

OpenStudy (welshfella):

the parametric equations for a circle with the centre at the origin sre y = sin t, x = cos t

ganeshie8 (ganeshie8):

check this http://www.mathopenref.com/coordparamcircle.html

OpenStudy (mpj4):

Thanks

ganeshie8 (ganeshie8):

You have \[(x-2)^2+y^2=4\] Notice that plugging in \(x=2+2\cos t\) makes the first term equal to \(4\cos^2 t\)

ganeshie8 (ganeshie8):

plugging in \(y = 2\sin t\) gives \[4\cos^2t + 4\sin^2t = 4\]

ganeshie8 (ganeshie8):

then use the identity \(\cos^2t+\sin^2 t = 1\)

OpenStudy (mpj4):

ugh... thanks anyway. You lost me on the last part there. It's ok, you don't need to bother explaining. I'll be closing this.